I just purchased a swimming pool with the dimensions of 15 feet across and 42 inches deep. Yes it is a circle. I need to know how many gallons of water it will hold. I am having a tanker brought out and need to know how much they need to bring because they charge by the gallon to bring it out.
Thank You,
William Beauchamp

Answer by oberobic(2304)

1 cubic foot of water is about 7.48 gallons.
.
The volume = Area of the circle * height of the column. This is the formula for the volume of a cylinder, which is what you will have.
.
Area = pi*r^2 = pi*7.5^2 = 176.71 square feet
.
Assuming you want to fill the pool to the rim, that is 42 inches, which = 3.5 ft.
.
3.5 ft * 176.71 ft^2 = 618.485 ft^3
.
618.485 ft^3 * 7.48 gallons/ft^3 = 4626.27 gallons
.
Done